# Deep Learning for Protein-Ligand Docking: Are We There Yet?

### v1

@@ Abstract
The effects of ligand binding on protein structures and their in vivo functions carry numerous implications for modern biomedical research and biotechnology development efforts such as drug discovery. Although several deep learning (DL) methods and benchmarks designed for protein-ligand docking have recently been introduced, to date no prior works have systematically studied the behavior of docking methods within the practical context of (1) predicted (apo) protein structures, (2) multiple ligands concurrently binding to a given target protein, and (3) having no prior knowledge of binding pockets. To enable a deeper understanding of docking methods’ real-world utility, we introduce PoseBench, the first comprehensive benchmark for practical protein-ligand docking. PoseBench enables researchers to rigorously and systematically evaluate DL docking methods for apo-to-holo protein-ligand docking and protein-ligand structure generation using both single and multi-ligand benchmark datasets, the latter of which we introduce for the first time to the DL community. Empirically, using PoseBench, we find that all recent DL docking methods but one fail to generalize to multi-ligand protein targets and also that template-based docking algorithms perform equally well or better for multi-ligand docking as recent single-ligand DL docking methods, suggesting areas of improvement for future work. Code, data, tutorials, and benchmark results are available at https://github.com/BioinfoMachineLearning/PoseBench.
@@ 1 Introduction
  • In contrast to several recent works on protein-ligand docking (Buttenschoen et al., 2024; Corso et al., 2024), the benchmark results we present in this work are all within the context of apo (i.e., predicted) protein structures without known binding pockets, which notably enhances the practicality and real-world utility of this study’s findings.
@@ 3.1 Preprocessed datasets
PoseBench provides users with three datasets with which to evaluate existing or new protein-ligand structure generation methods, the Astex Diverse and PoseBusters Benchmark datasets previously curated by Buttenschoen et al. (2024) as well as the CASP15 protein-ligand interaction (PLI) dataset that we have manually curated in this work.
@@
Likewise to Astex Diverse, for the PoseBusters Benchmark set, we used ESMFold to predict the apo complex structures of each of its proteins. After filtering out 28 complexes for which the corresponding protein structure could not be predicted on an 80GB A100 GPU, we RMSD-aligned the remaining 280 predicted protein structures while optimally weighting each complex’s protein-ligand interface in the alignment.
CASP15 dataset.
To assess the multi-ligand modeling capabilities of recent methods for protein-ligand structure generation, in this work, we introduce a curated version of the CASP15 PLI dataset introduced as a first-of-its-kind prediction category in the 15th Critical Assessment of Structure Prediction (CASP) competition (Robin et al., 2023) held in 2022. The CASP15 PLI set is originally comprised of 23 protein-ligand complexes, where we subsequently filter out 4 complexes based on (1) whether the CASP organizers ultimately assessed predictions for the complexes; (2) whether they are RNA-ligand complexes with no interacting protein chains; or (3) whether we could obtain a reasonably accurate prediction of the complex’s multimeric protein chains using either ESMFold or AlphaFold-based structure prediction on an 80GB A100 GPU (selecting for each complex the prediction which yielded the lowest-RMSD protein complex structure). Following this initial filtering step, we optimally align each remaining complex’s predicted protein structures to the corresponding ground-truth protein-(multi-)ligand structures, weighting each of the complex’s protein-ligand binding sites in the structural alignment.
@@
Sequence identity overlap. Note that for all three of the test datasets described above and listed in Table 1, we do not perform an analysis of the sequence identity overlap between these test datasets’ proteins and those of e.g., the PDBBind (time-split) training dataset, as (1) not all DL-based docking methods use PDBBind as their respective training datasets and (2) leaving the test complexes unfiltered according to sequence identity should, in principle, reflect many real-world use cases of these methods in which several (new) protein targets they are presented with may or may not be similar to what the methods have "seen" during training. Nevertheless, for an investigation of the sequence identity overlap between e.g., the PoseBusters Benchmark set and PDBBind, we refer interested readers to Buttenschoen et al. (2024).
@@ 4 Methods and Experimental Setup
As representative algorithms for conventional protein-ligand docking, we include AutoDock Vina (v1.2.5) (Trott and Olson, 2010) as well as a template-based modeling method for accurate ligand-protein complex structure prediction (TULIP) that we introduce in this work. To represent predictive ML docking algorithms, we include FABind (Pei et al., 2024) as well as the recently released version of RoseTTAFold 2 for all-atom structural modeling (i.e., RoseTTAFold-All-Atom) (Krishna et al., 2024). Lastly, for generative ML docking algorithms, we include DynamicBind (Lu et al., 2024), NeuralPLexer (Qiao et al., 2024), and the latest version of DiffDock referred to as DiffDock-L (Corso et al., 2024) which is designed with binding site generalization as a key aim.
Additionally, we provide a method ensembling baseline (Ensemble) that uses (multi-)ligand structural consensus ranking (Con) (Roy et al., 2023) to rank its ligand structure predictions selected from the (intrinsically method-ranked) top-40 ligand conformations produced by each conventional, predictive, and generative ML method. This ensembling baseline is included for each of PoseBench’s docking tasks to answer the question, "Which method produces the most consistent conformations in interaction with a protein complex?".
@@
Formats for generative methods are as follows: \2 DiffDock-L is provided with a predicted protein structure and (fragment) ligand SMILES strings and is then tasked with producing (multiple rank-ordered) ligand conformations (for each fragment) for the given protein. Note that DiffDock-L does not natively support multi-ligand SMILES string inputs, so in this work, we propose a modified inference procedure for DiffDock-L which autoregressively presents each (fragment) ligand SMILES string to the model while providing the same predicted protein structure to the model in each inference iteration. Notably, as an inference-time modification, this sampling formulation permits multi-ligand sampling yet cannot model multi-ligand interactions directly and therefore often produces ligand-ligand steric clashes. \2 As a single-ligand generative docking method, DynamicBind adopts the same input and output formats as DiffDock-L with the following exceptions: (1) the predicted input protein structure can be modified in response to (fragment) ligand docking, and (2) the autoregressive inference procedure we adapted from that of DiffDock-L now provides DynamicBind with its own most recently generated protein structure in each (fragment) ligand inference iteration, thereby providing the model with partial multi-ligand interaction context. Nonetheless, for both DiffDock-L and DynamicBind, such modified inference procedures highlight the importance in future work of retraining such generative methods directly on multi-ligand complexes to address such inference-time compromises. \2 Lastly, as a natively multi-ligand structure generation model pretrained using various 3D molecular and protein data sources, NeuralPLexer receives as its inputs a (multi-chain) protein sequence, a predicted protein (template) structure, as well as (fragment) ligand SMILES strings. The method is then tasked with producing (multiple rank-ordered) protein-ligand structure conformations for each input complex, using the method’s average predicted per-ligand heavy atom lDDT score (Mariani et al., 2013) for rank-ordering.
Prediction and evaluation procedures. With the prediction formats described above for each type of method, the protein-ligand complex structures each method produces are subsequently evaluated using various structural accuracy and molecule validity metrics depending on whether the targets are single or multi-ligand complexes.
@@
Multi-ligand evaluation. Following CASP15’s official scoring procedure for protein-ligand complexes (Robin et al., 2023), for multi-ligand targets, we report each method’s percentage of "correct" (binding site-superimposed) ligand conformations (RMSD $\leq$ 2 Å) as well as violin plots of the RMSD and PLI-specific lDDT scores of its protein-ligand conformations across all (fragment) ligands within the benchmark’s multi-ligand complexes (see Appendix C for these plots). Notably, this final metric, referred to lDDT-PLI, allows one to evaluate specifically how well each method can model protein-ligand structural interfaces. In the remainder of this work, we will describe and discuss the experimental results produced using our benchmark and their implications for the development of future protein-ligand structure generation methods.
@@ 5.1 Generalization to new binding pockets implies single-ligand docking performance
We begin our investigations by evaluating the performance of each baseline method for single-ligand docking using the Astex Diverse and PoseBusters Benchmark datasets. Notably, for results on the PoseBusters Benchmark dataset, we perform an additional analysis where we apply post-prediction (fixed-protein) relaxation to each method’s generated ligand conformations using molecular dynamics simulations (Eastman and Pande, 2010), as originally proposed by Buttenschoen et al. (2024). As shown in Figures 2 and 3, DiffDock-L achieves the best overall performance across the two datasets both with and without applying relaxation to its generated structures. Closely behind in performance for the PoseBusters Benchmark dataset are DynamicBind and RoseTTAFold-All-Atom following structural relaxation. Interestingly, without relaxation, AutoDock Vina combined with DiffDock-L’s predicted binding pockets achieves the second-best performance on the PoseBusters Benchmark dataset, which suggests that DiffDock-L is currently the only single-ligand deep learning method that presents a better intrinsic understanding of biomolecular physics for docking than conventional modeling tools. For interested readers, in Appendix C.1, we report additional results e.g., in terms of RMSD violin plots for both the Astex Diverse and PoseBusters Benchmark datasets.
@@ 5.2 Molecule pretraining implies multi-ligand docking performance
Another interesting observation is that TULIP (following structural relaxation) outperforms the docking success rates of single-ligand deep learning docking methods such as DiffDock-L and DynamicBind in the context of multi-ligand docking, which suggests room for future improvement in the multi-ligand modeling capabilities of these recent deep learning baselines. To further inspect each method’s understanding of biomolecular physics for docking, in Figure 5 we report each method’s percentage of predicted complexes (whether correct or not) for which all ligand conformations in the complex are jointly considered valid according to the PoseBusters software suite (i.e., PB-Valid). In short, in the context of multi-ligands, we find that AutoDock Vina followed by DiffDock-L are tied in terms of their PoseBusters validity rates following structural relaxation, with our ensembling consensus baseline (i.e., Ensemble (Con)) as well as DynamicBind shortly behind. As an addendum, we note that NeuralPLexer’s (DynamicBind’s) predictions seem to be frequently selected by Ensemble (Con) for single (multi)-ligand targets, which suggests that NeuralPLexer (DynamicBind) produces the most consistent (i.e., similar) ligand poses for a given single (multi)-ligand protein complex. For interested readers, in Appendix C.2, we report additional results e.g., in terms of lDDT-PLI and RMSD violin plots for both the total available CASP15 targets as well as the publicly available ones.
@@ Conclusions
Availability. The PoseBench codebase as well as tutorial notebooks are available under an MIT license at https://github.com/BioinfoMachineLearning/PoseBench. Preprocessed datasets and benchmark method predictions are also available, of which the Astex Diverse and PoseBusters Benchmark datasets are associated with a CC-BY 4.0 license and of which the CASP15 dataset, as a mixture of publicly and privately available resources, is partially licensed. In particular, 15 (4 single-ligand and 11 multi-ligand targets) of the 19 CASP15 protein-ligand complexes evaluated with PoseBench are publicly available, whereas the remaining 4 (2 single-ligand and 2 multi-ligand targets) are confidential and, for the purposes of future benchmarking and reproducibility, must be requested directly from the CASP organizers. Lastly, our use of the PoseBusters software suite for molecule validity checking is permitted under a BSD-3-Clause license.
@@
This work was supported by a U.S. NSF grant (DBI2308699) and two U.S. NIH grants (R01GM093123 and R01GM146340). Additionally, this work was performed using high performance computing infrastructure provided by Research Support Services at the University of Missouri-Columbia (DOI: 10.32469/10355/97710). We would like to thank Martin Buttenschoen and Andriy Kryshtafovych for graciously allowing us to use in Figure 1 the PoseBusters and CASP logos, respectively. Lastly, we would like to thank Arian Jamasb for providing insightful feedback on an early version of this manuscript.
@@
Appendix A Broader Impacts

### v2

@@ Abstract
The effects of ligand binding on protein structures and their in vivo functions carry numerous implications for modern biomedical research and biotechnology development efforts such as drug discovery. Although several deep learning (DL) methods and benchmarks designed for protein-ligand docking have recently been introduced, to date no prior works have systematically studied the behavior of docking methods within the practical context of (1) using predicted (apo) protein structures for docking (e.g., for broad applicability); (2) docking multiple ligands concurrently to a given target protein (e.g., for enzyme design); and (3) having no prior knowledge of binding pockets (e.g., for pocket generalization). To enable a deeper understanding of docking methods’ real-world utility, we introduce PoseBench, the first comprehensive benchmark for practical protein-ligand docking. PoseBench enables researchers to rigorously and systematically evaluate DL docking methods for apo-to-holo protein-ligand docking and protein-ligand structure generation using both single and multi-ligand benchmark datasets, the latter of which we introduce for the first time to the DL community. Empirically, using PoseBench, we find that all recent DL docking methods but one fail to generalize to multi-ligand protein targets and also that template-based docking algorithms perform equally well or better for multi-ligand docking as recent single-ligand DL docking methods, suggesting areas of improvement for future work. Code, data, tutorials, and benchmark results are available at https://github.com/BioinfoMachineLearning/PoseBench.
@@ 1 Introduction
  • In contrast to several recent works on protein-ligand docking (Buttenschoen et al., 2024; Corso et al., 2024a), the benchmark results we present in this work are all within the context of apo (i.e., predicted) protein structures without known binding pockets, which notably enhances the practicality and real-world utility of this study’s findings.
@@ 3.1 Preprocessed datasets
PoseBench provides users with four datasets with which to evaluate existing or new protein-ligand structure generation methods, the Astex Diverse and PoseBusters Benchmark (DockGen) datasets previously curated by Buttenschoen et al. (2024) ((Corso et al., 2024a)) as well as the CASP15 protein-ligand interaction (PLI) dataset that we have manually curated in this work.
@@
Likewise to Astex Diverse, for the PoseBusters Benchmark set, we used ESMFold to predict the apo complex structures of each of its proteins. After filtering out 28 complexes for which the corresponding protein structure could not be predicted on an 80GB A100 GPU, we RMSD-aligned the remaining 280 predicted protein structures while optimally weighting each complex’s protein-ligand interface in the alignment. For the DockGen dataset, we refer readers to Appendix F.1.
CASP15 dataset. To assess the multi-ligand modeling capabilities of recent methods for protein-ligand structure generation, in this work, we introduce a curated version of the CASP15 PLI dataset introduced as a first-of-its-kind prediction category in the 15th Critical Assessment of Structure Prediction (CASP) competition (Robin et al., 2023) held in 2022. The CASP15 PLI set is originally comprised of 23 protein-ligand complexes, where we subsequently filter out 4 complexes based on (1) whether the CASP organizers ultimately assessed predictions for the complexes; (2) whether they are RNA-ligand complexes with no interacting protein chains; or (3) whether we could obtain a reasonably accurate prediction of the complex’s multimeric protein chains using either ESMFold or AlphaFold-based structure prediction on an 80GB A100 GPU (selecting for each complex the prediction which yielded the lowest-RMSD protein complex structure). Following this initial filtering step, we optimally align each remaining complex’s predicted protein structures to the corresponding ground-truth protein-(multi-)ligand structures, weighting each of the complex’s protein-ligand binding sites in the structural alignment.
@@
Sequence identity overlap. Note that for all four of the test datasets described above and listed in Table 1, we do not perform an analysis of the sequence identity overlap between these test datasets’ proteins and those of e.g., the PDBBind (time-split) training dataset, as (1) not all DL-based docking methods use PDBBind as their respective training datasets and (2) leaving the test complexes unfiltered according to sequence identity should, in principle, reflect many real-world use cases of these methods in which several (new) protein targets they are presented with may or may not be similar to what the methods have "seen" during training. Nevertheless, for an investigation of the sequence identity overlap between e.g., the PoseBusters Benchmark set and PDBBind, we refer interested readers to Buttenschoen et al. (2024). Furthermore, in Appendix E, we analyze the different types and frequencies of protein-ligand interactions natively found within the Astex Diverse, PoseBusters Benchmark, DockGen, and CASP15 datasets, respectively, to quantify the diversity of the (predicted) interactions each dataset can be used to evaluate.
@@ 4 Methods and experimental setup
As representative algorithms for conventional protein-ligand docking, we include AutoDock Vina (v1.2.5) (Trott and Olson, 2010) as well as a template-based modeling method for accurate ligand-protein complex structure prediction (TULIP) that we introduce in this work. To represent predictive ML docking algorithms, we include FABind (Pei et al., 2024) as well as the recently released version of RoseTTAFold 2 for all-atom structural modeling (i.e., RoseTTAFold-All-Atom) (Krishna et al., 2024). Lastly, for generative ML docking algorithms, we include DynamicBind (Lu et al., 2024), NeuralPLexer (Qiao et al., 2024), and the latest version of DiffDock referred to as DiffDock-L (Corso et al., 2024a) which is designed with binding site generalization as a key aim. Notably, AlphaFold 3 (Abramson et al., 2024) does not support generic SMILES string inputs, so we cannot benchmark it.
Additionally, we provide a method ensembling baseline (Ensemble) that uses (multi-)ligand structural consensus ranking (Con) (Roy et al., 2023) to rank its ligand structure predictions selected from the (intrinsically method-ranked) top-40 ligand conformations produced by each conventional, predictive, and generative ML method. This ensembling baseline is included to answer the question, "Which method produces the most consistent conformations in interaction with a protein complex?".
@@
Formats for generative methods are as follows: \2 DiffDock-L is provided with a predicted protein structure and (fragment) ligand SMILES strings and is then tasked with producing (multiple rank-ordered) ligand conformations (for each fragment) for the given protein. Note that DiffDock-L does not natively support multi-ligand SMILES string inputs, so in this work, we propose a modified inference procedure for DiffDock-L which autoregressively presents each (fragment) ligand SMILES string to the model while providing the same predicted protein structure to the model in each inference iteration (reporting for each complex the average confidence score over all iterations). Notably, as an inference-time modification, this sampling formulation permits multi-ligand sampling yet cannot model multi-ligand interactions directly and therefore often produces ligand-ligand steric clashes. \2 As a single-ligand generative docking method, DynamicBind adopts the same input and output formats as DiffDock-L with the following exceptions: (1) the predicted input protein structure can be modified in response to (fragment) ligand docking; (2) the autoregressive inference procedure we adapted from that of DiffDock-L now provides DynamicBind with its own most recently generated protein structure in each (fragment) ligand inference iteration, thereby providing the model with partial multi-ligand interaction context; and (3) iteration-averaged confidence scores and predicted affinities are reported for each complex. Nonetheless, for both DiffDock-L and DynamicBind, such modified inference procedures highlight the importance in future work of retraining such generative methods directly on multi-ligand complexes to address such inference-time compromises. \2 Lastly, as a natively multi-ligand structure generation model pretrained using various 3D molecular and protein data sources, NeuralPLexer receives as its inputs a (multi-chain) protein sequence, a predicted protein (template) structure, as well as (fragment) ligand SMILES strings. The method is then tasked with producing (multiple rank-ordered) protein-ligand structure conformations for each input complex, using the method’s average predicted per-ligand heavy atom lDDT score (Mariani et al., 2013) for rank-ordering.
Prediction and evaluation procedures. Using the prediction formats above, the protein-ligand complex structures each method produces are subsequently evaluated using various structural accuracy and molecule validity metrics depending on whether the targets are single or multi-ligand complexes.
@@
Multi-ligand evaluation. Following CASP15’s official scoring procedure for protein-ligand complexes (Robin et al., 2023), for multi-ligand targets, we report each method’s percentage of "correct" (binding site-superimposed) ligand conformations (RMSD $\leq$ 2 Å) as well as violin plots of the RMSD and PLI-specific lDDT scores of its protein-ligand conformations across all (fragment) ligands within the benchmark’s multi-ligand complexes (see Appendix F for these plots). Notably, this final metric, referred to lDDT-PLI, allows one to evaluate specifically how well each method can model protein-ligand structural interfaces. In the remainder of this work, we will discuss our benchmark’s results and their implications for the development of future complex structure generation methods.
@@ 5.1 Generalization to new binding pockets implies single-ligand docking performance
We begin our investigations by evaluating the performance of each baseline method for single-ligand docking using the Astex Diverse and PoseBusters Benchmark datasets. Notably, for results on the PoseBusters Benchmark dataset, we perform an additional analysis where we apply post-prediction (fixed-protein) relaxation to each method’s generated ligand conformations using molecular dynamics simulations (Eastman and Pande, 2010), as originally proposed by Buttenschoen et al. (2024). As shown in Figures 2 and 3, DiffDock-L achieves the best overall performance across the two datasets both with and without applying relaxation to its generated structures. Closely behind in performance for the PoseBusters Benchmark dataset are DynamicBind and RoseTTAFold-All-Atom following structural relaxation. Interestingly, without relaxation, AutoDock Vina combined with DiffDock-L’s predicted binding pockets achieves the second-best performance on the PoseBusters Benchmark dataset, which suggests that DiffDock-L is currently the only single-ligand deep learning method that presents a better intrinsic understanding of biomolecular physics for docking than conventional modeling tools. For interested readers, in Appendix F.2, we report e.g., pocket-only PoseBusters Benchmark experiments and RMSD violin plots for both the Astex Diverse and PoseBusters Benchmark datasets.
@@ 5.2 Molecule pretraining implies multi-ligand docking performance
Another interesting observation is that TULIP (following structural relaxation) outperforms the docking success rates of single-ligand deep learning docking methods such as DiffDock-L and DynamicBind in the context of multi-ligand docking, which suggests room for future improvement in the multi-ligand modeling capabilities of these recent deep learning baselines. To further inspect each method’s understanding of biomolecular physics for docking, in Figure 5 we report each method’s percentage of predicted complexes (whether correct or not) for which all ligand conformations in the complex are jointly considered valid according to the PoseBusters software suite (i.e., PB-Valid). In short, in the context of multi-ligands, we find that AutoDock Vina followed by DiffDock-L are tied in terms of their PoseBusters validity rates following structural relaxation, with our ensembling consensus baseline (i.e., Ensemble (Con)) as well as DynamicBind shortly behind. As an addendum, we note that NeuralPLexer’s (DynamicBind’s) predictions seem to be frequently selected by Ensemble (Con) for single (multi)-ligand targets, which suggests that NeuralPLexer (DynamicBind) produces the most consistent (i.e., similar) ligand poses for a given single (multi)-ligand protein complex. For interested readers, in Appendix F.3, we report additional results e.g., in terms of lDDT-PLI and RMSD violin plots for both the total available CASP15 targets as well as the publicly available ones.
@@ Conclusions
Availability. The PoseBench codebase, documentation, and tutorial notebooks are available on GitHub under a permissive MIT license, with further licensing discussed in Appendix A.
@@
This work was supported by a U.S. NSF grant (DBI2308699) and two U.S. NIH grants (R01GM093123 and R01GM146340). Additionally, this work was performed using high performance computing infrastructure provided by Research Support Services at the University of Missouri-Columbia (DOI: 10.32469/10355/97710). We would like to thank Martin Buttenschoen and Andriy Kryshtafovych for graciously allowing us to use in Figure 1 the PoseBusters and CASP logos, respectively. We would also like to specifically thank Martin Buttenschoen for their assistance in running the PoseBusters software suite for various protein-ligand complexes. Lastly, we would like to thank Arian Jamasb for providing insightful feedback on an early version of this manuscript.
Checklist
  1. 1.
    For all authors…
    1. (a)
      Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope? [Yes] In this work, we have introduced PoseBench for comprehensive benchmarking of protein-ligand structure generation methods. The PoseBench benchmark results we present in this manuscript subsequently outline important research directions for future work on deep learning for protein-ligand docking including the importance of pretraining methods on large molecule corpora and evaluating them directly on multi-ligand docking targets.
    2. (b)
      Did you describe the limitations of your work? [Yes] We discuss the limitations of this study in Section Conclusions.
    3. (c)
      Did you discuss any potential negative societal impacts of your work? [Yes] In Appendix B, we discuss the broader impacts of this work as they relate to the benefits and potential risks of accurate protein-ligand structure generation methods.
    4. (d)
      Have you read the ethics review guidelines and ensured that your paper conforms to them? [Yes] Our newly proposed protein-ligand docking benchmark adheres to the ethics review guidelines to the best of our knowledge.
  2. 2.
    If you are including theoretical results…
    1. (a)
      Did you state the full set of assumptions of all theoretical results? [N/A]
    2. (b)
      Did you include complete proofs of all theoretical results? [N/A]
  3. 3.
    If you ran experiments (e.g. for benchmarks)…
    1. (a)
      Did you include the code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL)? [Yes] For full reproducibility and extensibility, all associated source code, data, tutorials, documentation, and benchmark method predictions are made freely available at https://github.com/BioinfoMachineLearning/PoseBench.
    2. (b)
      Did you specify all the training details (e.g., data splits, hyperparameters, how they were chosen)? [Yes] In Sections 3 and 4, we carefully list the dataset preparation and method inference steps we took to produce the results we report in Section 5. Additionally, the source code and data we have made freely available will enhance readers’ ability to understand our results.
    3. (c)
      Did you report error bars (e.g., with respect to the random seed after running experiments multiple times)? [Yes] For all generative baseline methods included in our experiments, we report the mean and standard deviation (across three independent runs) of each of the method’s corresponding performance metrics. We plot these statistical terms for each method using Matplotlib’s bar chart function [Hunter, 2007], specifically its height and yerr arguments, respectively. For completeness, like many previous works on benchmarking biomolecular methods in machine learning [Jamasb et al., 2024], we assume normally distributed errors when constructing our result plots.
    4. (d)
      Did you include the total amount of compute and the type of resources used (e.g., type of GPUs, internal cluster, or cloud provider)? [Yes] In Appendix C, we provide comprehensive details regarding the compute resources we utilized to run the experiments presented in this work.
  4. 4.
    If you are using existing assets (e.g., code, data, models) or curating/releasing new assets…
    1. (a)
      If your work uses existing assets, did you cite the creators? [Yes] Please see Appendix A.
    2. (b)
      Did you mention the license of the assets? [Yes] Please see Appendix A.
    3. (c)
      Did you include any new assets either in the supplemental material or as a URL? [Yes] As mentioned in Appendix A, we have made freely available on Zenodo [Morehead et al., 2024] curated, deep learning-friendly versions of the Astex Diverse, PoseBusters Benchmark, DockGen, and CASP15 protein-ligand interaction datasets for apo-to-holo protein-(multi-)ligand structure generation.
    4. (d)
      Did you discuss whether and how consent was obtained from people whose data you’re using/curating? [Yes] In Appendix A, we discuss the licensing and usage of each dataset that our benchmark curates in this work. In addition, we have received permission from the creators of the PoseBusters software suite and the CASP15 organizers to reference their work in our study.
    5. (e)
      Did you discuss whether the data you are using/curating contains personally identifiable information or offensive content? [N/A] The data curated in this study is of macromolecular nature, without any personally identifiable information associated.
  5. 5.
    If you used crowdsourcing or conducted research with human subjects…
    1. (a)
      Did you include the full text of instructions given to participants and screenshots, if applicable? [N/A]
    2. (b)
      Did you describe any potential participant risks, with links to Institutional Review Board (IRB) approvals, if applicable? [N/A]
    3. (c)
      Did you include the estimated hourly wage paid to participants and the total amount spent on participant compensation? [N/A]
[…]
[appendix]
@@
Appendix A Availability
The PoseBench codebase and tutorial notebooks are available under an MIT license at https://github.com/BioinfoMachineLearning/PoseBench. Preprocessed datasets and benchmark method predictions are available on Zenodo [Morehead et al., 2024] under a CC-BY 4.0 license, of which the Astex Diverse and PoseBusters Benchmark datasets [Buttenschoen et al., 2024] are associated with a CC-BY 4.0 license; of which the DockGen dataset [Corso et al., 2024a] is available under an MIT license; and of which the CASP15 dataset [Robin et al., 2023], as a mixture of publicly and privately available resources, is partially licensed. In particular, 15 (4 single-ligand and 11 multi-ligand targets) of the 19 CASP15 protein-ligand complexes evaluated with PoseBench are publicly available, whereas the remaining 4 (2 single-ligand and 2 multi-ligand targets) are confidential and, for the purposes of future benchmarking and reproducibility, must be requested directly from the CASP organizers. Lastly, our use of the PoseBusters software suite for molecule validity checking is permitted under a BSD-3-Clause license.
Appendix B Broader impacts